# Intelligent Framework for Human-Robot Collaboration: Safety, Dynamic Ergonomics, and Adaptive Decision-Making

### v1

@@ Abstract
The integration of collaborative robots into industrial environments has improved productivity, but has also highlighted significant challenges related to operator safety and ergonomics. This paper proposes an innovative framework that integrates advanced visual perception technologies, real-time ergonomic monitoring, and Behaviour Tree (BT)-based adaptive decision-making. Unlike traditional methods, which often operate in isolation or statically, our approach combines deep learning models (YOLO11 and SlowOnly), advanced tracking (Unscented Kalman Filter) and dynamic ergonomic assessments (OWAS), offering a modular, scalable and adaptive system. Experimental results show that the framework outperforms previous methods in several aspects: accuracy in detecting postures and actions, adaptivity in managing human-robot interactions, and ability to reduce ergonomic risk through timely robotic interventions. In particular, the visual perception module showed superiority over YOLOv9 and YOLOv8, while real-time ergonomic monitoring eliminated the limitations of static analysis. Adaptive role management, made possible by the Behaviour Tree, provided greater responsiveness than rule-based systems, making the framework suitable for complex industrial scenarios. Our system demonstrated a 92.5% accuracy in grasping intention recognition and successfully classified ergonomic risks with real-time responsiveness (average latency of 0.57 seconds), enabling timely robotic
@@
Human-robot collaboration, real-time ergonomics, adaptive decision-making, visual perception, Behaviour Trees, YOLO.
@@
TABLE I: Comparison of COCO Dataset Architects. The performance of the examined models is compiled in the table, which further emphasizes how YOLO11 uniquely combines accuracy, speed, and segmentation capacity. This combination presents it as the ideal approach to handle the operational difficulties of collaborative environments, therefore enhancing both efficiency and safety.
@@ I Introduction
Industrial automation has revolutionized manufacturing processes and enabled the integration of collaborative robotic systems, allowing robots and human operators to work together and enhancing the productivity and adaptability of production lines. This paradigm, termed Human-Robot Collaboration (HRC), has shown considerable productivity enhancements while also presenting intricate issues concerning safety, ergonomics, and flexibility in dinamics operational situations [1].
Figure 1: Schematic depiction of the proposed structure, illustrating the main modules and their interaction dynamics.
The prevention of Work-Related Musculoskeletal Disorders (WMSD) [2], which can arise due to incorrect postures or repetitive strain, is one of the most relevant issues. These disorders not only compromise the health of operators, but also negatively affect company productivity, increasing operating costs. As demonstrated by extensive studies on muscle fatigue in work environments [3], these issues require comprehensive monitoring solutions that can adapt to dynamic working conditions and provide timely interventions. Traditional technologies such as OWAS (Ovako Working Posture Analysis System) and REBA (Rapid Entire Body Assessment) offer useful tools for identifying ergonomic risks, but are inadequate for the continuous and dynamic monitoring required in today’s complex industrial environments.
To the best of our knowledge, no existing framework in the literature synergistically integrates advanced technologies such as visual detection, real-time ergonomic monitoring, and adaptive decision-making through Behavior Trees (BTs) for collaborative industrial settings. Current approaches address these challenges in isolation or only partially, limiting their practical applicability [4]. This work aims to bridge this gap by proposing an innovative framework that combines state-of-the-art technologies to enhance safety, ergonomics, and efficiency in industrial environments.
The proposed framework, illustrated in Fig. 1, stands out for its non-invasive nature and its ability to adapt to complex and dynamic operational scenarios. It integrates advanced visual detection technologies (YOLOv11 and OpenPose) for posture recognition, a modular decision-making system based on Behavior Trees (BTs) for dynamic human-robot role adaptation, and continuous ergonomic assessment methods to prevent physical risk situations. This integration approach addresses the safety challenges highlighted in industrial collaboration studies [1], while providing the adaptability required for varied manufacturing tasks and the non-invasiveness that traditional sensor-based systems [5] often lack. This unified approach enables real-time monitoring and optimization of human-robot interactions, enhancing both safety and overall productivity.
@@ II-A Visual Perception and Object Detection
The visual detection and segmentation of objects are crucial to ensure safety and efficiency in human-robot interactions [6]. In our framework, we chose to adopt YOLO11 [7], a state-of-the-art model that offers an excellent balance between accuracy, speed of inference and native segmentation capabilities, as reported in Table I.
Compared to its predecessors, YOLO11 [7] stands out due to significant optimisations in speed and accuracy. Models such as YOLOv8 [8], YOLOv9 [9] and YOLOv10 [10], while representing advances over previous generations, show limitations in handling complex scenarios or operational speed. At the same time, EfficientDet [11], although appreciated for its scalability and energy efficiency, is less competitive in dynamic contexts due to higher inference times and lack of native segmentation.
Even when compared to more complex models such as Mask R-CNN [12] and Faster R-CNN [13, 14], YOLO11 [7] stands out due to its lighter computational load, which makes it ideal for real-time applications where speed is essential to ensure effective responses. These features make it particularly suitable for complex collaborative scenarios, where rapid movements and frequent overlaps require highly responsive and accurate technology.
Integrated into our framework, YOLO11[7] demonstrated a remarkable ability to detect postures and objects even in scenarios characterised by complex movements, frequent overlaps and variable geometries. Its accuracy and speed proved essential to improve the tracking and monitoring of interactions, allowing the system to quickly adapt to operational requirements.
In conclusion, as is clear from Table I, YOLO11[7] represents a significant evolution from the original YOLO concept [6] and a significant step ahead of the alternatives due to its ability to combine high accuracy, accurate segmentation and operational speed. This evolution addresses key limitations identified in early object detection systems while maintaining the real-time performance critical for human-robot collaborative environments [1], where even milliseconds of delay can impact safety and operational efficiency. This combination makes it an ideal tool to address the challenges of human-robot collaboration, improving both safety and operational efficiency.
@@ II-B Real-Time Tracking
Real-time tracking is crucial to ensure safety and fluidity in human-robot interactions. The Speed and Separation Monitoring (SSM) method proposed by Marvel et al. [15] monitors the speed and distance between human and robot to prevent collisions, but the lack of integration with advanced visual perception technologies limits its fluidity in complex environments . In contrast, Simple Online and Realtime Tracking (SORT) [16] offers a more responsive solution, using Kalman filtering [17] for rapid detection coupling, but suffers from reduced performance in the presence of occlusions or non-linear movements .
In our framework, the integration of YOLO11 with the Unscented Kalman Filter (UKF) [18] improves the handling of non-linear dynamics, overcoming the limitations of previous approaches such as SORT. This approach builds upon fundamental filtering principles [17] while extending their applicability to the complex, non-linear movements typical of human operators in industrial settings. This integration offers more reliable tracking in complex and rapidly changing industrial environments, , making operations safer and more efficient as emphasized by safety guidelines in collaborative robotic systems [1].
@@ II-D Dynamic Role Allocation with Behaviour Trees
Dynamic role allocation is crucial for efficient collaboration, especially in industrial settings where conditions can change rapidly. Rozo et al. [26] have employed probabilistic models to adapt roles based on human demonstrations, but these models struggle to handle dynamic industrial environments.
In our framework, we use Behaviour Trees (BT) [27], which offer a more modular and flexible decision-making structure, adapting robotic tasks to operational conditions detected in real time. Recent advances in unified architectures for dynamic role allocation and collaborative task planning in mixed human-robot teams [28] further demonstrate the importance of adaptable frameworks. Merlo et al. [29] proposed a similar framework based on ergonomic indicators, aimed at reducing the risk of musculoskeletal fatigue through intelligent role allocation. Our work extends this approach by integrating continuous monitoring of operators’ physical conditions to improve overall safety, addressing key concerns highlighted in comprehensive surveys of HRC methodologies [23] and providing a more responsive system for dynamic task environments.
@@ II-E Real-Time Ergonomic Assessment
Our framework overcomes these limitations by integrating dynamic ergonomic analysis with advanced computer vision technologies such as OpenPose, enabling continuous monitoring of postures in real time [36]. This real-time analysis capability addresses a fundamental gap identified in traditional ergonomic evaluation methods [30, 33], which typically require manual observation and cannot adapt to rapidly changing work conditions. This approach not only prevents injuries related to incorrect posturesas cataloged in traditional ergonomic assessment methods [31, 32], but also enables an immediate adaptive response, improving safety and reducing operator muscle fatigue through interventions that align with established ergonomic principles [34].
Figure 2: Overview of the system’s perception and decision-making pipeline, showing the interplay between human pose tracking, action recognition, object detection, and ergonomic fatigue assessment.
@@ II-G Proposed Framework Improvements
  • Dynamic Ergonomic Evaluation: Provides continuous and real-time evaluation, improving injury prevention due to incorrect postures or repetitive movements. This continuous monitoring represents a significant advancement over the static approaches criticized in previous studies [35] and addresses key concerns related to musculoskeletal disorders [3], with our experimental results showing effective classification of ergonomic risks in real-time (average latency of 0.57 seconds).
  • Advanced Action Perception: The integration of OpenPose with SlowOnly allows for an in-depth understanding of human actions, identifying movements that could lead to excessive physical stress. This integration addresses the challenges of temporal recognition identified in prior work [24] and extends the capabilities of standard pose estimation systems [19], achieving 95.83% accuracy in action recognition during our validation tests.
  • Intelligent Automation: The behaviour tree-based decision-making system allows selective and adaptive intervention by the robot, optimising operational efficiency without interfering with human activities. Our behavior tree implementation extends existing role allocation frameworks [29] with enhanced adaptability, while addressing safety concerns highlighted in comprehensive HRC surveys [23, 4], demonstrating rapid response times (0.07 seconds on average) to critical ergonomic conditions.
@@ III Proposed Framework Architecture
The proposed framework, in Fig. 2, combines advanced computer vision techniques, human action recognition, and ergonomic evaluation, enabling synergy between human operator and robot in collaborative environments. This architecture enables real-time monitoring of operators’ operating conditions, with an adaptive decision-making system managing dynamic task allocation between human and robot via a Behaviour Tree (BT).
@@
III-A Visual Perception Module
The Visual Perception Module is responsible for detecting, tracking, and calculating the physical characteristics of objects and operators’ hands. Data flow from YOLO11 and tracking via Unscented Kalman Filter (UKF) provides parallel inputs that feed into the decision-making system and ergonomic evaluation.
III-B YOLO11: Object Detection
For object detection and the generation of accurate segmentation masks, the YOLO11 model was used, which provides detailed data on the position and size of detected objects, which is essential for ensuring safety in human-robot collaboration environments. The output of the model is represented by a vector $\mathbf{y}$, which includes position, size, confidence and segmentation of objects:
@@ III-B YOLO11: Object Detection
Where $\lambda_{text{conf}}$, $\lambda_{text{cls}}$ and $\lambda_{text{seg}}$ are scalar weights that balance the relative importance of each component of the loss.
This formulation allows the system to detect and segment objects in complex scenarios, ensuring accurate separation even in the presence of overlaps or irregular shapes.
III-C Calculation of Area, Volume, and Volumetric Weight: Input from YOLO11
Using the data provided by YOLO11, the system performs the parallel calculation of the area, volume, and volumetric weight of objects, which are key inputs for ergonomic evaluation using OWAS.
Area:
@@ III-C Calculation of Area, Volume, and Volumetric Weight: Input from YOLO11
Volumetric Weight:
Once the volume is estimated, the volumetric weight is calculated by dividing the volume by a standard conversion factor:
$P_{\text{volumetric}}=\frac{V_{\text{object}}}{\text{conversion factor}}$
@@
Figure 3: OWAS evaluation and result interpretation [39].
III-C1 Kalman Unscented Filter (UKF)
In parallel with calculating the physical properties of objects, the
III-C2 Unscented Kalman Filter
(UKF) is used to track the positions and movements of objects and operators’ hands in real time. UKF is particularly well suited to handle nonlinear dynamics typical of complex movements in industrial environments.
Tracking is handled through a state vector $\mathbf{x}$ that includes the positions, velocities, and orientations of objects and hands:
@@ III-C2 Unscented Kalman Filter
Here, $\mathbf{p}$ denotes the position in 3D space, $\mathbf{v}$ the linear velocity in 3D space, and $\mathbf{q}$ the orientation expressed in quaternions to avoid the ambiguities that arise in the use of Euler angles.
The state update is done by a transition function $\mathbf{f(x)}$ that predicts the evolution of the variables over time. This function takes into account the linear and rotational dynamics of the system and is expressed as follows:
@@
  • $\mathbf{p}$ is the updated position as a function of linear velocity $\mathbf{v}$ and time $dt$,
  • $\mathbf{v}$ represents velocity, held constant within the update interval $dt$. This choice is justified by the high update rate of the system, which makes speed changes between two consecutive updates negligible,
  • $\mathbf{q}$ represents the updated orientation using the quaternion product with the angular velocity $\boldsymbol{\omega}$ and time $dt$.
@@
$\hat{\mathbf{x}}_{k+1}=f(\mathbf{x}_{k})$
@@
After estimating the future state with the transition function, the UKF compares this estimate with actual observations obtained from depth sensors, such as the RealSense D435 camera. The measurement function $\mathbf{h(x)}$maps the estimated state onto observables, i.e. the position and orientation of objects and hands:
@@
The measurement predicted at time $k$,
$\hat{\mathbf{z}}_{k}$
is obtained by applying the measurement function to the state predicted at time $k$:
@@
Subsequently, the new actual observations ${\mathbf{z}}_{k}$ are compared with those predicted estimates. The error between the predicted estimate $\hat{\mathbf{z}}_{k}$ and the actual observations $\mathbf{z}_{k}$ is used to update and correct the estimated state, via the Kalman gain $K$.The final correction of the state $\mathbf{x}_{k+1}$, is made with the following equation:$k$:
@@
In this way, the estimated state is updated using new sensory observations, improving accuracy in real time.
@@
The stability and accuracy of the system strongly depend on a proper definition of the covariance matrices. The process noise covariance matrix $\mathbf{Q}$ represents the uncertainty in the system model, while the observation noise matrix $\mathbf{R}$ handles the uncertainty associated with sensor measurements:
$\mathbf{Q}=\text{diag}([100,\dots,100])\quad\text{e}\quad\mathbf{R}=\text{diag}([0.01,\dots,0.01])$
The initial state covariance matrix $\mathbf{P}$, defined as a scaled identity matrix, also plays a key role in the accuracy of the first estimates:
@@
The correct choice of these parameters is essential to ensure that the UKF system maintains stability and accuracy even in the presence of uncertainties in the process or measurements, minimizing estimation error through the entire operational cycle.
III-D Human Actions Recognition Module
The
III-E Human Actions Recognition Module
integrates
III-F OpenPose
for pose detection and
III-G SlowOnly
for large time-scale action recognition.
III-H OpenPose: Pose Detection
OpenPose detects the joints of the human body in 3D using a perspective projectionciteiodice2022learning. Each i-th joint is represented by the coordinates $(x,y,z)$ and a confidence score $c$:

### v2

# Intelligent Framework for Human-Robot Collaboration: Dynamic Ergonomics and Adaptive Decision-Making

@@ Abstract
The integration of collaborative robots into industrial environments has improved productivity, but has also highlighted significant challenges related to operator safety and ergonomics. This paper proposes an innovative framework that integrates advanced visual perception, continuous ergonomic monitoring, and adaptive Behaviour Tree decision-making to overcome the limitations of traditional methods that typically operate as isolated components. Our approach synthesizes deep learning models, advanced tracking algorithms, and dynamic ergonomic assessments into a modular, scalable, and adaptive system. Experimental validation demonstrates the framework’s superiority over existing solutions across multiple dimensions: the visual perception module outperformed previous detection models with 72.4% mAP@50:95; the system achieved high accuracy in recognizing operator intentions (92.5%); it promptly classified ergonomic risks with minimal latency (0.57 seconds); and it dynamically managed robotic interventions with exceptionally responsive decision-making capabilities (0.07 seconds), representing a 56% improvement over benchmark systems. This comprehensive solution provides a robust platform for enhancing human-robot collaboration in industrial environments by prioritizing ergonomic safety, operational efficiency, and real-time adaptability.
@@
Human-robot collaboration, Real-time ergonomics, Visual perception, Adaptive decision-making, Intention recognition, Integrated safety framework
@@
TABLE I: Performance comparison of object detection models on COCO dataset. The table presents model complexity (Parameters in millions, FLOPs in billions), detection accuracy (mean Average Precision at IoU thresholds 0.5-0.95), processing speed (inference latency in milliseconds), and segmentation capabilities. YOLO models show progressive improvements across generations, with YOLO11-L achieving 53.4% mAP at 6.20ms latency while maintaining segmentation functionality, demonstrating the trade-offs between computational requirements and performance metrics relevant for real-time applications.
@@ I Introduction
Industrial automation has revolutionized manufacturing processes and enabled the integration of collaborative robotic systems, allowing robots and human operators to work together and enhancing the productivity and adaptability of production lines. This paradigm, termed Human-Robot Collaboration (HRC), has shown considerable productivity enhancements while also presenting intricate issues concerning safety, ergonomics, and flexibility in dynamic operational situations [1].
Figure 1: Proposed HRC framework integrating object detection and tracking (YOLO11, Unscented Kalman Filter), 3D human pose tracking (OpenPose), action recognition (SlowOnly), OWAS-based ergonomic assessment, and a Behaviour Trees decision-making system. Solid arrows represent direct data flows between primary functional components, while dashed arrows indicate indirect or complementary information exchanges that support the main processing pipeline. The architecture facilitates non-invasive continuous monitoring, dynamic human-robot role allocation, and ergonomic-based interventions to enhance safety, ergonomics, and productivity in industrial environments.
The prevention of Work-Related Musculoskeletal Disorders (WMSD) [2], which can arise due to incorrect postures or repetitive strain, is one of the most relevant issues. These disorders not only compromise the health of operators, but also negatively affect company productivity, increasing operating costs. As demonstrated by extensive studies on muscle fatigue in work environments [3], these issues require comprehensive monitoring solutions that can adapt to dynamic working conditions and provide timely interventions. Traditional methodologies such as the Ovako Working Posture Analysis System (OWAS) and Rapid Entire Body Assessment (REBA) offer useful tools for identifying ergonomic risks, but are inadequate for the continuous monitoring required in today’s complex industrial environments. To the best of our knowledge, no existing framework in the literature synergistically integrates advanced technologies such as visual detection, continuous ergonomic monitoring, and adaptive decision-making through Behaviour Trees (BT) for collaborative industrial settings. Current approaches address these challenges in isolation or only partially, limiting their practical applicability [4]. This work aims to bridge this gap by proposing an innovative framework that combines state-of-the-art technologies to enhance safety, ergonomics, and efficiency in industrial environments. The proposed framework, illustrated in Fig. 1, stands out for its non-invasive nature and its ability to adapt to complex and dynamic operational scenarios. It integrates advanced visual detection technologies (YOLO11 with Unscented Kalman Filter tracking and physical properties calculation, and OpenPose with 3D perspective projection) for comprehensive posture and object recognition, an action recognition module using temporal analysis for movement pattern interpretation, a modular decision-making system based on BTs with conditional and composite nodes for dynamic human-robot role allocation and ergonomic-based intervention, and continuous OWAS-based ergonomic assessment methods to prevent WMSD and other physical risk situations. This integration approach addresses the safety challenges highlighted in industrial collaboration studies [1], while providing the adaptability required for varied manufacturing tasks and the non-invasiveness that traditional sensor-based systems [5] often lack. This unified approach enables continuous monitoring and optimization of human-robot interactions, enhancing both safety and overall productivity.
@@ II-A Visual Perception and Object Detection
Visual perception constitutes the foundational layer of human-robot collaborative systems, determining the quality of all subsequent processing stages. The evolution of object detection methodologies reveals a critical transition from accuracy-prioritizing two-stage detectors to efficiency-oriented single-stage architectures designed for time-sensitive applications. Early two-stage detection methods, such as R-CNN [6] and Fast R-CNN [7], achieved high accuracy but were computationally inefficient. Faster R-CNN [7] introduced region proposal networks to reduce computational requirements; however, their latency remains relatively high, limiting real-time application potential in industrial contexts (see Table I).
Single-stage detection frameworks, initiated by the YOLO architecture [8], significantly reduced latency by formulating detection as a regression problem. Subsequent iterations addressed specific limitations: YOLOv2 [9] improved localization accuracy, YOLOv3 [10] enhanced multi-scale detection, and YOLOv4 [11] optimized performance across platforms. More recently, YOLOv8 incorporated segmentation capabilities (YOLO-seg), enabling more detailed scene understanding critical for industrial tasks. YOLOv9 introduced advanced attention mechanisms improving feature representation, whereas YOLOv10 prioritized latency optimization, compromising segmentation capabilities. YOLO11 balanced these trade-offs, offering enhanced segmentation performance and low latency, making it highly suitable for real-time HRC applications, as highlighted by metrics in Table I.
Alternative architectures like EfficientDet [12] explored compound scaling to efficiently balance model complexity and detection accuracy, achieving good speed but lacking native segmentation capabilities. Mask R-CNN [13] extended Faster R-CNN by incorporating instance segmentation, providing high accuracy but at significant computational costs and latency unsuitable for real-time scenarios (Table I).
Detection performance degradation under real-world industrial conditions such as occlusions and variable lighting was demonstrated by Dodge and Karam [14], and Michaelis et al. [15]. YOLO11 addresses these limitations effectively, as it maintains high performance in variable industrial environments, which our experimental validation confirmed.
@@ II-B Real-Time Tracking
Real-time tracking is crucial to ensure safety and fluidity in human-robot interactions. The Speed and Separation Monitoring (SSM) method proposed by Marvel et al. [16] monitors the speed and distance between human and robot to prevent collisions, but the lack of integration with advanced visual perception technologies limits its fluidity in complex environments . In contrast, Simple Online and Realtime Tracking (SORT) [17] offers a more responsive solution, using Kalman filtering [18] for rapid detection coupling, but suffers from reduced performance in the presence of occlusions or non-linear movements, which are common challenges in dynamic industrial environments.
More recent developments, such as DeepSORT and ByteTrack [19], improve robustness under occlusion conditions. Our proposed approach integrates YOLO11 with an Unscented Kalman Filter (UKF) [20], extending traditional Kalman filter principles to effectively manage complex human movements typical of industrial environments.
@@ II-D Dynamic Role Allocation with Behaviour Trees
Dynamic role allocation is essential for efficient collaboration, especially in industrial settings characterized by rapidly changing conditions. Traditional methods, such as probabilistic models based on human demonstrations proposed by Rozo et al. [28], often exhibit limitations in adapting flexibly to dynamic environments. Similarly, Finite State Machines (FSMs) face scalability challenges due to rapidly increasing complexity [29], and Markov Decision Processes (MDPs) [30] often require extensive computational resources, limiting their practical application in real-time adaptive scenarios.
To overcome these limitations, our framework employs Behaviour Trees (BT) [31], which provide a modular, hierarchical, and adaptable decision-making structure. Behaviour Trees facilitate rapid adaptation of robotic tasks by continuously monitoring operational conditions. Compared to probabilistic models, FSMs, and MDPs, BTs significantly simplify the integration of multiple real-time input streams, enhancing responsiveness and adaptability.
Recent advancements in unified architectures for dynamic role allocation [32] underline the importance of such adaptable frameworks. Merlo et al. [33] have also proposed ergonomic-driven role allocation aimed at reducing musculoskeletal fatigue. Our work further develops this concept by integrating continuous ergonomic monitoring, significantly improving overall safety, flexibility, and operational responsiveness in dynamically evolving industrial environments [25].
@@ II-E Real-Time Ergonomic Assessment
Our approach overcomes these limitations by integrating dynamic ergonomic analysis with advanced computer vision technologies such as OpenPose, enabling continuous monitoring of postures in real time [40]. This real-time analysis capability addresses a fundamental gap identified in traditional ergonomic evaluation methods [34, 37], which typically require manual observation and cannot adapt to rapidly changing work conditions. This approach not only prevents injuries related to incorrect postures as cataloged in traditional ergonomic assessment methods [35, 36], but also enables an immediate adaptive response, improving safety and reducing operator muscle fatigue through interventions that align with established ergonomic principles [38].
Figure 2: Overview of the proposed human-robot collaboration framework. The architecture integrates three input modules (3D Human Pose Tracking, Action Recognition, and Object Detection and Tracking) that feed into the OWAS ergonomic assessment system. The ergonomic evaluation classifies postures according to back, arms, legs, and load parameters, incorporating a temporal sliding window to track sustained postures. Based on this assessment, the Behaviour Tree manages dynamic role allocation between human operator and robot, prioritizing robot intervention when biomechanical fatigue is detected. This integrated approach ensures operational efficiency while preserving worker ergonomic safety.
@@ II-G Proposed Framework Improvements
  • Continuous Ergonomic Monitoring: Real-time ergonomic assessments improve proactive injury prevention by dynamically identifying high-risk postures and movements.
  • Enhanced Action Recognition: Integration of OpenPose 3D and SlowOnly enables accurate detection of subtle and repetitive actions indicative of fatigue or ergonomic risks.
  • Intelligent Task Adaptation: Utilizing Behaviour Trees facilitates adaptable robot interventions based on real-time conditions, optimizing safety and efficiency.
@@ III Proposed Framework Architecture
The proposed framework, illustrated in Fig. 2, combines advanced computer vision techniques, human action recognition, and ergonomic evaluation, enabling synergy between human operator and robot in collaborative environments. This architecture enables continuous monitoring of operators’ physical conditions, with an adaptive decision-making system managing dynamic task allocation between human and robot via a BT.
@@
III-A Object Detection and Tracking Module
This module is responsible for detecting, tracking, and calculating the physical characteristics of objects in the environment, providing essential data for both the decision-making system and ergonomic evaluation.
III-A1 Object Detection with YOLO11
For object detection and the generation of accurate segmentation masks, the YOLO11 model is used, which provides detailed data on the position and size of detected objects, essential for ensuring safety in human-robot collaboration environments. The output of the model is represented by a vector $\mathbf{y}$, which includes position, size, confidence and segmentation of objects:
@@ III-A1 Object Detection with YOLO11
Where:
  • $\mathcal{L}_{\text{box}}$ represents the bounding box regression loss that penalizes localization errors in the predicted object boundaries
  • $\mathcal{L}_{\text{conf}}$ represents the confidence loss that quantifies the model’s certainty about the presence of an object
  • $\mathcal{L}_{\text{cls}}$ represents the classification loss that penalizes incorrect category predictions
  • $\mathcal{L}_{\text{seg}}$ represents the segmentation loss that measures the accuracy of the predicted binary mask
  • $\lambda_{\text{conf}}$, $\lambda_{\text{cls}}$ and $\lambda_{\text{seg}}$ are scalar weights that balance the relative importance of each component of the loss
This multi-objective loss function enables joint optimization of spatial localization and semantic understanding. This formulation allows the system to detect and segment objects in complex scenarios, ensuring accurate separation even in the presence of overlaps or irregular shapes.
III-A2 Physical Property Calculation for Ergonomic Assessment
Using the data provided by YOLO11, the system performs the parallel calculation of the area, volume, and volumetric weight of objects, which are key inputs for ergonomic evaluation using the OWAS methodology.
Area
@@ III-A2 Physical Property Calculation for Ergonomic Assessment
Volumetric Weight
Once the volume is estimated, the volumetric weight, $P_{\text{vol}}$, is calculated by dividing the volume by a standard conversion factor:
$P_{\text{vol}}=\frac{V_{\text{object}}}{\text{conversion factor}}$
@@
III-A3 Unscented Kalman Filter for Movement Tracking
In parallel with calculating the physical properties of objects, the UKF is used to track the positions and movements of objects and operators’ hands in real time. The UKF is particularly suitable for handling nonlinear dynamics typical of complex movements in industrial environments.
Tracking is handled through a state vector $\mathbf{x}\in\mathbb{R}^{30}$ that includes the positions, velocities, and orientations of objects and hands:
@@ III-A3 Unscented Kalman Filter for Movement Tracking
Here, $\mathbf{p}\in\mathbb{R}^{3}$ denotes the position in 3D space, $\mathbf{v}\in\mathbb{R}^{3}$ the linear velocity in 3D space, and $\mathbf{q}\in\mathbb{R}^{4}$ the orientation expressed in quaternions to avoid the ambiguities that arise in the use of Euler angles.
The state update is conducted through a transition function $\mathbf{f}:\mathbb{R}^{30}\rightarrow\mathbb{R}^{30}$ that predicts the evolution of the variables over time. This function takes into account the linear and rotational dynamics of the system and is expressed as follows:
@@
  • $\mathbf{p}\in\mathbb{R}^{3}$ is the updated position as a function of linear velocity $\mathbf{v}$ and time interval $dt\in\mathbb{R}^{+}$,
  • $\mathbf{v}\in\mathbb{R}^{3}$ represents velocity, held constant within the update interval $dt$. This choice is justified by the high update rate of the system (30 Hz), which makes speed changes between two consecutive updates negligible,
  • $\mathbf{q}\in\mathbb{R}^{4}$ represents the updated orientation using the quaternion product $\otimes$ with the angular velocity $\boldsymbol{\omega}\in\mathbb{R}^{3}$ and time interval $dt$,
  • $\mathbf{\Omega}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{4}$ is a function that converts angular velocity into quaternion derivative format.
@@
$\hat{\mathbf{x}}_{k+1}=\mathbf{f(\mathbf{x}_{k})}$
where $\hat{\mathbf{x}}_{k+1}\in\mathbb{R}^{30}$ represents the predicted state at time step $k+1$, and $\mathbf{x}_{k}\in\mathbb{R}^{30}$ is the current state at time step $k$.
@@
After estimating the future state with the transition function, the UKF compares this estimate with actual observations obtained from depth sensors, such as the RealSense D435 camera. The measurement function $\mathbf{h}:\mathbb{R}^{30}\rightarrow\mathbb{R}^{21}$ maps the estimated state onto observables, i.e., the position and orientation of objects and hands:
@@
The measurement vector $\mathbf{z}_{k}\in\mathbb{R}^{21}$ consists of the direct observations of positions and orientations, excluding velocities which are not directly measurable. The measurement predicted at time $k$, $\hat{\mathbf{z}}_{k}\in\mathbb{R}^{21}$ is obtained by applying the measurement function to the state predicted at time $k$:
@@
Subsequently, the new actual observations ${\mathbf{z}}_{k}$ are compared with those predicted estimates. The error between the predicted estimate $\hat{\mathbf{z}}_{k}$ and the actual observations $\mathbf{z}_{k}$ is used to update and correct the estimated state, via the Kalman gain $K\in\mathbb{R}^{30\times 21}$. The final correction of the state $\mathbf{x}_{k+1}$ is made with the following equation:
@@
The Kalman gain $K$ is computed internally by the UKF algorithm and optimally weights the relative importance of the model prediction versus the new measurements, based on their respective uncertainties.
@@
The stability and accuracy of the system strongly depend on a proper definition of the covariance matrices. The process noise covariance matrix $\mathbf{Q}\in\mathbb{R}^{30\times 30}$ represents the uncertainty in the system model, while the observation noise matrix $\mathbf{R}\in\mathbb{R}^{21\times 21}$ handles the uncertainty associated with sensor measurements:
$\mathbf{Q}=\text{diag}([100,\dots,100])\quad\text{and}\quad\mathbf{R}=\text{diag}([0.01,\dots,0.01])$
The initial state covariance matrix $\mathbf{P}\in\mathbb{R}^{30\times 30}$, defined as a scaled identity matrix, also plays a key role in the accuracy of the first estimates:
@@
where $\mathbf{I}_{30}\in\mathbb{R}^{30\times 30}$ is the identity matrix of dimension 30.
The UKF implementation uses the Merwe scaled sigma point parameterization with $\alpha=10^{-3}$, $\beta=2.0$, and $\kappa=0$, which determines how the sigma points are distributed around the mean to capture the system’s nonlinearities. To ensure numerical stability, a small regularization term ($1e^{-8}$) is added to the covariance matrices, and positive definiteness is enforced through eigenvalue decomposition when necessary.
The correct selection of these parameters is essential to ensure that the UKF system maintains stability and accuracy even in the presence of uncertainties in the process or measurements, minimizing estimation error throughout the entire operational cycle.
Figure 3: OWAS (Ovako Working Posture Analysis System) evaluation matrix. The left section displays the coding system for back (1-4), arms (1-3), legs (1-7), and load (1-3) postures. The right section presents the action category matrix with color-coded risk levels (green: no action required; yellow: corrective actions required in the near future; orange: corrective actions should be done as soon as possible; red: corrective actions for improvement required immediately) based on combinations of back, arms, legs, and load values.
III-B 3D Human Pose Tracking Module
The 3D Human Pose Tracking Module is responsible for detecting and analyzing the operator’s posture in real-time, providing crucial data for ergonomic evaluation.
III-B1 Pose Detection with OpenPose
OpenPose detects the joints of the human body in 3D using a perspective projection [44]. Each i-th joint is represented by the coordinates $(x,y,z)$ and a confidence score $c$: